A fruit drop from the top of tree 20m tall the Tim take the fruit to reach the ground


1
Expert's answer
2021-02-22T16:03:39-0500
"h=\dfrac{1}{2}gt^2,""t=\sqrt{\dfrac{2h}{g}}=\sqrt{\dfrac{2\cdot20\ m}{9.8\ \dfrac{m}{s^2}}}=2.02\ s."